OpenStudy (anonymous):

Integrate. sqrt(1+x^2)/x

OpenStudy (anonymous):

I think it should be \[\int\limits_{}^{} 1/\sin(x)\cos^2x\]

OpenStudy (mimi_x3):

Or start from \[\int\frac{\sqrt{1+x^{2}}}{x} dx\]

OpenStudy (anonymous):

You could started from there, I got up to 1/sinxcos^2x

OpenStudy (mimi_x3):

Assuming that you did the other part correct then.. \[\int \frac{1}{sinxcos^2{x}}\] wait..are you sure about that though?

OpenStudy (lgbasallote):

 \[\sec \theta = \sqrt{1+ x^2}\] \[\tan \theta = x\] \[\sec^2 \theta d\theta = dx\] \[\int \frac{\sqrt{1+x^2}}{x}dx \implies \frac{\sec \theta(\sec^2 \theta d\theta )}{\tan \theta}\]

OpenStudy (anonymous):

I'm sure that \[\int\limits_{}^{} \sec^3x/\tan(x)\]

OpenStudy (anonymous):

That turns into 1/cos^2xsinx

OpenStudy (lgbasallote):

it does?

OpenStudy (lgbasallote):

i mean yeah it does

OpenStudy (anonymous):

I know I can't do u sub.

OpenStudy (lgbasallote):

there's only one thing you can do...integration by parts

OpenStudy (lgbasallote):

\[\int \frac{1}{\sin x \cos^2 x}dx \implies \int \csc x \sec^2 x dx\]

OpenStudy (anonymous):

But can I integrate that?

OpenStudy (lgbasallote):

well i was just spitting out ideas....

OpenStudy (anonymous):

That's why I didn't change it to that because I knew there was nothing to do there.

OpenStudy (lgbasallote):

well i know i can make 1 = sin^2 x + cos^2 x

OpenStudy (lgbasallote):

\[\int \frac{1}{\sin x \cos^2 x}dx \implies \int \frac{ \sin^2 x + \cos^2 x}{\sin x \cos^2 x}dx\]

OpenStudy (anonymous):

OH! That's good

OpenStudy (lgbasallote):

\[\int \frac{\sin^2 x}{\sin x \cos ^2 x}dx + \int \frac{\cos^2 x}{\sin x \cos^2 x}dx\]

OpenStudy (lgbasallote):

yup saw the answer now :D i'll let you do the good stuff ;)

OpenStudy (anonymous):

Ah, yes, yes, I see it too.

OpenStudy (lgbasallote):

nice! congrats